OpenStudy (pokemon23):

I need help to review my geometry exam that is tomorrow @ganeshie8

OpenStudy (pokemon23):

Identify each statement and determine the truth value. 1) 1+1 -2 or 3 less then 1 2) Halloween is not a season. 2) "If today is Monday, and then tomorrow is Tuesday." Write the converse, inverse,contrapositive.

ganeshie8 (ganeshie8):

logic !

OpenStudy (pokemon23):

I failed my first exam with a 64.....

ganeshie8 (ganeshie8):

oh dont wry i see u have very good concepts, you will get 100 this time :)

OpenStudy (pokemon23):

I placed a bet with my friend

ganeshie8 (ganeshie8):

logic is very easy IF we remember few things :- start wid below, 1) OR 2) AND

OpenStudy (pokemon23):

or- disjunction

ganeshie8 (ganeshie8):

if you see OR between two statements. test if the left statement is true test if the right statement is true if ATLEAST ONE statement is true, then the overall statement is true

ganeshie8 (ganeshie8):

1) 1+1 -2 or 3 less then 1

ganeshie8 (ganeshie8):

1) 1+1 = 2 or 3 less then 1 its like that right ?

OpenStudy (pokemon23):

correct

ganeshie8 (ganeshie8):

carry out the tests

ganeshie8 (ganeshie8):

1) 1+1 = 2 or 3 less then 1 left side statement :- ``` 1+1 = 2 ``` right side statemetn :- ``` 3 < 1 ```

OpenStudy (pokemon23):

yeah

ganeshie8 (ganeshie8):

Is left side true ? Is right side true ?

OpenStudy (pokemon23):

left side is true and right side is false

ganeshie8 (ganeshie8):

1) 1+1 = 2 or 3 less then 1 left side statement :- ``` 1+1 = 2 TRUE ``` right side statemetn :- ``` 3 < 1 FALSE ```

ganeshie8 (ganeshie8):

Since ATLEAST ONE statement is TRUE, the overall given statement is TRUE.

ganeshie8 (ganeshie8):

btw, OR = disjunction

OpenStudy (pokemon23):

:D thanks

ganeshie8 (ganeshie8):

we simply need to test left/right side statements conclude

ganeshie8 (ganeshie8):

2) Halloween is not a season.

ganeshie8 (ganeshie8):

NOT = negation

OpenStudy (pokemon23):

so the statement is true

ganeshie8 (ganeshie8):

yup ! its a NEGATION, and the truth value is TRUE

OpenStudy (pokemon23):

right

OpenStudy (pokemon23):

the next part I get confused

ganeshie8 (ganeshie8):

its very easy, and that q will come in exam for sure tomorrow

OpenStudy (pokemon23):

by next part I mean to say next question

OpenStudy (pokemon23):

I like that ;)

OpenStudy (pokemon23):

No you deleted o_O

ganeshie8 (ganeshie8):

remember this :- given statement :- \(\large \text{if a, then b}\) converse :- ``` if b, then a ``` inverse :- ``` if NOT a, then NOT b ``` contrapositive :- ``` if NOT b, then NOT a ```

ganeshie8 (ganeshie8):

"If today is Monday, then tomorrow is Tuesday." Write the converse, inverse,contrapositive. a = If today is Monday b = tomorrow is Tuesday

ganeshie8 (ganeshie8):

try writing converse/inverse/contrapostive

OpenStudy (pokemon23):

Ok for converse If Tomorrow is is Tuesday, then today is Monday.

ganeshie8 (ganeshie8):

Correct

OpenStudy (pokemon23):

For inverse its If today is not Monday, then tomorrow is not Tuesday.

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

OpenStudy (pokemon23):

For Contrapostive If tomorrow is not Tuesday, then today is Not Tuesday.

ganeshie8 (ganeshie8):

thats it !

OpenStudy (pokemon23):

can you tell me what a conjunction is?

OpenStudy (pokemon23):

Thank you for the help ganeshie8

ganeshie8 (ganeshie8):

AND = conjunction

ganeshie8 (ganeshie8):

its the sister of disjunction, but a very strict sister lol it requires both left/right to be TRUE

OpenStudy (pokemon23):

alright

ganeshie8 (ganeshie8):

1) 1+1 = 2 AND 3 less then 1

ganeshie8 (ganeshie8):

take that

ganeshie8 (ganeshie8):

we already evaluated left side = TRUE right side = FALSE right ?

OpenStudy (pokemon23):

yeah

ganeshie8 (ganeshie8):

so conjunction of those (AND) wud be FALSE, cuz AND requires both to be TRUE

OpenStudy (pokemon23):

yep

ganeshie8 (ganeshie8):

conjunction(AND) :- BOTH left/right have to be true, for AND to be true. disjunction(OR) :- ATLEAST ONE of left/right have to be true, for OR to be true.

ganeshie8 (ganeshie8):

do few more examples, u wil get confidence :)

OpenStudy (pokemon23):

alright :) thank you very much :3

ganeshie8 (ganeshie8):

np :) all the best for tomorrow exam !